# Protect Before Generate: Error Correcting Codes within Discrete Deep Generative Models

### v1

@@ C.3 Generation
Model BER WER
uncoded 5 0.002 0.008
coded 5/50 0.007 0.034
coded 5/80 0.004 0.021
coded 5/100 0.009 0.045
uncoded 8 0.015 0.071
coded 8/80 0.020 0.147
coded 8/160 0.021 0.160
coded 8/240 0.023 0.167
uncoded 10 0.057 0.373
coded 10/100 0.030 0.258
coded 10/200 0.034 0.282
coded 10/300 0.041 0.331
Table 4: Evaluation of the log-likelihood (LL) in MNIST.
@@
Appendix D CIFAR10 results
@@ D.2 Reconstruction and generation
In Table 6 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 23, where the coded models better capture the details in the images. We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for the explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
@@
Table 5: Evaluation of reconstruction performance in CIFAR10 with different model configurations.
Model PSNR (train) PSNR (test)
uncoded 70 17.985 17.596
coded 70/700 23.790 17.731
coded 70/1400 24.555 18.008
coded 70/2100 25.551 18.401
uncoded 100 18.509 18.334
coded 100/1000 24.754 18.229
coded 100/2000 24.866 18.927
coded 100/3000 25.646 18.920
uncoded 130 18.951 18.758
coded 130/1300 25.007 18.887
coded 130/2600 25.460 19.416
coded 130/3900 $\mathbf{25.515}$ $\mathbf{19.292}$
@@
Model BER WER FID
uncoded 70 0.162 1.000 177.524
coded 70/700 0.101 1.000 104.977
coded 70/1400 0.088 0.999 104.078
coded 70/2100 0.090 0.999 102.795
uncoded 100 0.182 1.000 172.063
coded 100/1000 0.123 1.000 107.887
coded 100/2000 0.114 1.000 101.182
coded 100/3000 0.138 1.000 107.287
uncoded 130 0.197 1.000 164.138
coded 130/1300 0.144 1.000 109.905
coded 130/2600 0.164 1.000 110.250
coded 130/3900 0.185 1.000 108.561
Appendix E Tiny ImageNet results
@@ E.2 Reconstruction and generation
In Table 8 we quantify the quality of the reconstructions measuring the PSNR in both training and test sets. Coded models yield higher PSNR values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 28, where the coded models better capture the details in the images. We observe that the coded model yields images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for the explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
@@
Table 7: Evaluation of reconstruction performance in Tiny ImageNet with different model configurations.
Model PSNR (train) PSNR (test)
uncoded 70 15.598 15.402
coded 70/700 18.156 15.158
coded 70/1400 18.396 15.419
coded 70/2100 18.228 15.789
uncoded 100 16.012 15.774
coded 100/1000 18.298 15.677
coded 100/2000 18.647 15.892
coded 100/3000 18.729 16.167
uncoded 130 16.278 16.009
coded 130/1300 18.719 15.901
coded 130/2600 18.818 $\mathbf{16.329}$
coded 130/3900 $\mathbf{19.020}$ 16.288
@@
Model BER WER FID
uncoded 70 0.143 1.000 265.474
coded 70/700 0.096 0.998 171.993
coded 70/1400 0.104 1.000 170.496
coded 70/2100 0.096 0.998 176.245
uncoded 100 0.164 1.000 234.358
coded 100/1000 0.099 1.000 153.743
coded 100/2000 0.097 1.000 162.889
coded 100/3000 0.098 1.000 163.049
uncoded 130 0.200 1.000 219.003
coded 130/1300 0.129 1.000 165.064
coded 130/2600 0.114 1.000 164.759
coded 130/3900 0.128 1.000 170.603
We also evaluate the model in the image generation task. In Fig. 29, we show examples of randomly generated images using different model configurations in Tiny ImageNet. These results are consistent with the ones obtained in reconstruction, as we observe that the coded models can generate more detailed and diverse images. Additionally, we obtained the FID score using the test set and 10k generated samples. For this, we used the implementation available at https://github.com/mseitzer/pytorch-fid. We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.
Appendix F IWAE results
@@
One could draw a parallel between the coded DVAE with repetition codes and the well-known IWAE [3], but the two approaches are fundamentally different. In the IWAE, independent samples are drawn from the variational posterior and propagated independently through the generative model to obtain a tighter variational bound on the marginal log-likelihood. In our method, we jointly propagate the output of the ECC encoder through the generative model, obtaining a single prediction and exploiting the introduced known correlations in the variational approximation of the posterior. In the case of repetition codes, the ECC encoder outputs are repeated bits, or repeated probabilities in the case of soft encoding. However, our approach extends beyond repetition codes, opening a new field for improved inference in discrete latent variable models.
In this work, we specifically utilize the redundancy introduced by the repetition code to correct potential errors made by the encoder through a soft decoding approach, leading to a more accurate approximation of $p(\boldsymbol{m}|\boldsymbol{x})$ and an improved proposal for sampling. The results obtained in the coded DVAE case cannot be achieved by training the uncoded DVAE with the IWAE objective. The following results compare the uncoded IWAE model with the coded DVAE trained on FMNIST. While the uncoded model shows slight performance gains with an increasing number of IWAE samples (which improves the evidence lower bound), it still underperforms compared to the coded model. Furthermore, when using 20 and 30 IWAE samples, the metrics slightly declined compared to using 10 samples, likely due to overfitting, as we applied a common early stopping point.
@@
Table 9: Comparison of the metrics obtained with our method and the uncoded DVAE trained with the IWAE objective.
Model BER WER Entropy Acc. Conf. Acc. PSNR
uncoded 8 0.089 0.384 0.467 0.594 0.595 15.598
uncoded 8 IWAE 10 samples 0.063 0.372 1.309 0.617 0.640 14.282
uncoded 8 IWAE 20 samples 0.075 0.447 1.391 0.634 0.651 14.237
uncoded 8 IWAE 30 samples 0.074 0.438 1.564 0.619 0.641 13.757
coded 8/80 0.021 0.144 2.905 0.750 0.816 17.318
coded 8/160 0.027 0.189 3.637 0.783 0.831 17.713
coded 8/240 0.037 0.231 4.000 0.799 0.893 17.861
Appendix G Hierarchical coded DVAE results
Inspired by polar codes [2], we present a hierarchical coded DVAE with two layers of latent bits, as illustrated in Figure 31. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Concurrently, the bits in the second layer, $\boldsymbol{m}_{2}$, are linearly combined with $\boldsymbol{m}_{1}$ following $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$, considering a binary field or Galois field. The resulting vector is then encoded with another repetition code to produce $\boldsymbol{c}_{2}$, which is subsequently modulated into $\boldsymbol{z}_{2}$. Finally, both $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ are concatenated and passed through the decoder network to generate $\boldsymbol{x}$. The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model. Inference follows a similar approach to the one employed in the coded DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
@@ Appendix G Hierarchical coded DVAE results
We adopt the same variational family as used in the standard Coded VAE; however, in this case, we incorporate both hierarchical levels, leading to
@@
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ is calculated following the same approach as in the coded DVAE with repetition codes, computing the all-are-zero and all-are-ones products of probabilities of the bits in $\boldsymbol{c}_{1}$ that are copies of the same message bit. The posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ considering both the encoder’s output and the inferred posterior distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$. Note that in this case, the decoder outputs the probabilities for both $\boldsymbol{c}_{1}$ and $\boldsymbol{c}_{2}$, with $\boldsymbol{c}_{2}$ being the encoded version of the linear combination $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$. Consequently, we first obtain $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ following the same approach as in the coded DVAE with repetition codes, and determine $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ as
$q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(p_{u}),$ (13)
$p_{u}=q_{\boldsymbol{\eta}}(m_{1,2,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,2,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x}).$ (14)
@@
$q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})=\prod_{u=1}^{M}\text{Ber}(q_{u}),$ (15)
$q_{u}=q_{\boldsymbol{\eta}}(m_{1,u}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,u}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,u}=1|\boldsymbol{x}).$ (16)
Next, we apply the soft encoding approach to incorporate the repetition code structure at both levels of the hierarchy.The posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ are repeated to obtain $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{1}|\boldsymbol{x})$, and the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ are repeated to produce $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{2}|\boldsymbol{x})$. Utilizing the reparameterization trick from Eq. 5, we sample $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$, concatenate them to form $\boldsymbol{z}$, and pass this through the decoder to generate $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$. The model is trained by maximizing the ELBO, given by
@@
We obtained results on the FMNIST dataset using a model with 5 information bits per branch and repetition rates of $R=1/10$ and $R=1/20$. In this case, we applied the same code rate to both branches, although varying code rates could be used to control the level of protection at each hierarchy level. Tables 10 and 11 present the metrics obtained for the different configurations. Specifically, Table 10 shows the overall metrics obtained with this structure, and Table 11 compares the error metrics across the two hierarchy levels. As expected, $\mathbf{m}_{2}$ shows poorer error metrics compared to $\mathbf{m}_{1}$, since the model provides more redundancy to $\mathbf{m}_{1}$ incorporating it in both branches. Although the overall metrics and generation quality are somewhat similar to those of the coded DVAE with 10 information bits (see tables in Figures 2 and 4), the introduced hierarchy results in a more interpretable latent space. In this setup, $\mathbf{m}_{1}$ captures global features (such as clothing types in the FMNIST dataset), while $\mathbf{m}_{2}$ controls individual features, as we can observe in Figure 32, where we show examples of the model’s generative outputs for fixed $\mathbf{m}_{1}$ and random samples of $\mathbf{m}_{2}$.
Figure 32: Examples of generated images using the hierarchical coded DVAE with (a) a 5/50 repetition code in each branch, and (b) a 5/100 repetition code in each branch. In all the examples provided, $\boldsymbol{m}_{1}$ was fixed while $\boldsymbol{m}_{2}$ was randomly sampled.
Table 10: Comparison of the obtained metrics for the coded DVAE with polar codes with different configurations, which we refer to as ’hierarchical coded DVAE’.
@@
Appendix H Ablation study
@@
H.1 Ablation study on the hyperparameter $\beta$
@@
H.1.1 Coded model
We first evaluate the influence of the parameter $\beta$ in coded models. We take as a reference the coded model with 8 information bits and a rate $R=1/30$, and train it using $\beta=5,10,15,20$. We assess the performance of the model in reconstruction and generation tasks. We observe the model is fairly robust, achieving similar performance across configurations in most metrics.
@@ H.1.2 Uncoded model
H.2 Ablation study on the number of trainable parameters
A consistent architecture was employed across all experiments, which is detailed in Section A. However, since the introduction of the code alters the dimensionality of the latent space, it is necessary to adjust the encoder’s output and the decoder’s input. This results in an augmentation of the trainable parameters in the coded cases compared to their uncoded counterparts.
@@ H.2.1 Evaluation
Appendix I Evaluating log-likelihood using the soft-encoding model

### v2

@@ E.3 Generation
Model BER BER MAP WER WER MAP
uncoded 5 0.002 0.001 0.008 0.004
coded 5/50 0.007 0.000 0.034 0.000
coded 5/80 0.004 0.000 0.021 0.000
coded 5/100 0.009 0.000 0.045 0.000
uncoded 8 0.015 0.013 0.071 0.057
[…]
uncoded 10 0.057 0.052 0.373 0.353
[…]
Table 4: Evaluation of the LL in MNIST.
@@
Appendix F CIFAR10 results
@@ F.2 Reconstruction and generation
In Table 6 we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets. Coded models yield higher PSNR and SSIM values in train, and similar values in test, although the coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 25, where the coded models better capture the details in the images. We observe that the coded model produces images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that, to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
@@
Table 5: Reconstruction metrics in CIFAR10 with different model configurations.
Model
PSNR
SSIM
PSNR
SSIM
uncoded 70 17.985 0.340 17.596 0.324
coded 70/700 23.790 0.709 17.731 0.323
coded 70/1400 24.555 0.748 18.008 0.357
coded 70/2100 25.551 0.748 18.401 0.385
uncoded 100 18.509 0.381 18.334 0.370
coded 100/1000 24.754 0.754 18.229 0.375
coded 100/2000 24.866 0.761 18.927 0.432
coded 100/3000 25.646 0.793 18.920 0.426
uncoded 130 18.951 0.419 18.758 .408
coded 130/1300 25.007 0.767 18.887 0.426
coded 130/2600 25.460 0.784 $\mathbf{19.416}$ $\mathbf{0.464}$
coded 130/3900 $\mathbf{25.515}$ $\mathbf{0.785}$ 19.292 0.456
@@
Model BER
BER
MAP
WER
WER
MAP
FID
uncoded 70 0.162 0.162 1.000 0.999 177.524
coded 70/700 0.101 0.081 1.000 0.988 104.977
coded 70/1400 0.088 0.044 0.999 0.917 104.078
coded 70/2100 0.090 0.029 0.999 0.811 102.795
uncoded 100 0.182 0.179 1.000 1.000 172.063
coded 100/1000 0.123 0.104 1.000 0.999 107.887
coded 100/2000 0.114 0.060 1.000 0.990 101.182
coded 100/3000 0.138 0.077 1.000 0.997 107.287
uncoded 130 0.197 0.194 1.000 1.000 164.138
coded 130/1300 0.144 0.115 1.000 0.999 109.905
coded 130/2600 0.164 0.102 1.000 1.000 110.250
coded 130/3900 0.185 0.132 1.000 1.000 108.561
Appendix G Tiny ImageNet results
@@ G.2 Reconstruction and generation
In Table 8, we quantify the quality of the reconstructions measuring the PSNR and the SSIM in both training and test sets. Coded models yield higher PSNR and SSIM values in train, and similar values in test, although coded models with lower rates outperform the rest of the configurations. However, the improvement in reconstruction is evident through visual inspection of Fig. 30, where the coded models better capture the details in the images. We observe that the coded model produces images that better resemble the structure of the dataset, while the uncoded DVAE cannot decouple spatial information from the images and project it in the latent space.
We hypothesize that, to adequately model complex images, transitioning to a hierarchical structure may be necessary. This would allow for the explicit modeling of both global and local information. However, despite employing this rather simple model, we observe that coded configurations outperform their uncoded counterparts in capturing colors and textures.
We also evaluate the model in the image generation task. In Fig. 31, we show examples of randomly generated images using different model configurations in Tiny ImageNet. These results are consistent with the ones obtained in reconstruction, where we observe that the coded models can generate more detailed and diverse images. Additionally, we obtained the FID score using the test set and 10k generated samples. For this, we used the implementation available at https://github.com/mseitzer/pytorch-fid. We can observe that the coded models significantly reduced the FID score in all the cases compared to their uncoded counterparts.
@@
Table 7: Reconstruction metrics in Tiny ImageNet with different model configurations.
Model
PSNR
SSIM
PSNR
SSIM
uncoded 70 15.598 0.207 15.402 0.196
coded 70/700 18.156 0.367 15.158 0.191
coded 70/1400 18.396 0.383 15.419 0.203
coded 70/2100 18.228 0.369 15.789 0.223
uncoded 100 16.012 0.229 15.774 0.215
coded 100/1000 18.298 0.378 15.677 0.218
coded 100/2000 18.647 0.404 15.892 0.232
coded 100/3000 18.729 0.408 16.167 0.248
uncoded 130 16.278 0.243 16.009 0.228
coded 130/1300 18.719 0.409 15.900 0.233
coded 130/2600 18.818 $\mathbf{0.433}$ $\mathbf{16.329}$ $\mathbf{0.259}$
coded 130/3900 $\mathbf{19.020}$ 0.415 16.288 $\mathbf{0.259}$
@@
Model BER
BER
MAP
WER
WER
MAP
FID
uncoded 70 0.143 0.140 1.000 0.999 265.474
coded 70/700 0.096 0.072 0.998 0.978 171.993
coded 70/1400 0.104 0.066 1.000 0.972 170.496
coded 70/2100 0.096 0.034 0.998 0.832 176.245
uncoded 100 0.164 0.162 1.000 1.000 234.358
coded 100/1000 0.099 0.074 1.000 0.996 153.743
coded 100/2000 0.097 0.053 1.000 0.981 162.889
coded 100/3000 0.098 0.035 1.000 0.925 163.049
uncoded 130 0.200 0.198 1.000 1.000 219.003
coded 130/1300 0.129 0.107 1.000 0.999 165.064
coded 130/2600 0.114 0.060 1.000 0.996 164.759
coded 130/3900 0.128 0.070 1.000 0.999 170.603
Appendix H IWAE results
@@
One could draw a parallel between the Coded-DVAE with repetition codes and the well-known IWAE [Burda et al., 2016], but the two approaches are fundamentally different. In the IWAE, independent samples are drawn from the variational posterior and propagated independently through the generative model to obtain a tighter variational bound on the marginal log-likelihood. In our method, we jointly propagate the output of the ECC encoder through the generative model, obtaining a single prediction and exploiting the introduced known correlations in the variational approximation of the posterior. In the case of repetition codes, the ECC encoder outputs are repeated bits, or repeated probabilities in the case of soft encoding. However, our approach extends beyond repetition codes, opening a new field for improved inference in discrete latent variable models.
In our approach, we specifically utilize the redundancy introduced by the repetition code to correct potential errors made by the encoder through a soft decoding approach. This results in a more accurate approximation of the posterior $p(\boldsymbol{m}|\boldsymbol{x})$ and an improved proposal for sampling, resulting in improved performance over uncoded models, even those trained with the tighter IWAE bound. To compute the IWAE objective in the uncoded case, we draw samples from the posterior using the reparameterization trick described in Eq. (5), and compute the importance weights as $\boldsymbol{w}=\frac{p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})p(\boldsymbol{z})}{q_{\boldsymbol{\eta}}(\boldsymbol{z}|\boldsymbol{x})}$. Here, the prior and posterior over $\boldsymbol{z}$ are obtained by marginalizing out $\boldsymbol{m}$ in $p(\boldsymbol{z},\boldsymbol{m})=p(\boldsymbol{m})p(\boldsymbol{z}|\boldsymbol{m})$ and $q_{\boldsymbol{\eta}}(\boldsymbol{m},\boldsymbol{z}|\boldsymbol{x})=q_{\boldsymbol{\eta}}(\boldsymbol{m}|\boldsymbol{x})p(\boldsymbol{z}|\boldsymbol{m})$, respectively.
Additionally, we implemented two extensions of the IWAE bound to confirm that the observed improvements are not due to a known shortcoming of the reparametrized gradient estimator of the IWAE bound. Following Daudel et al. [2023], we implemented the VR-IWAE generalization, which introduces a hyperparameter $\alpha\in[0,1)$ into the objective function. This formulation interpolates between the IWAE bound (recovered when $\alpha=0$) and the ELBO (recovered when $\alpha=1$). For our experiments, we selected a midpoint value of $\alpha=0.5$. Additionally, we implemented the doubly-reparametrized (DR) gradient estimator associated with the VR-IWAE objective, as described in Daudel et al. [2023]. We trained the models on FMNIST using 10, 20, and 30 posterior samples, matching the number of repetitions used in our coded models. We included an additional experiment training an uncoded model with the IWAE bound considering 100 samples.
From the results, outlined in Table 9, we observe that training uncoded models with tighter bounds (even if we use bounds that are meant to become tighter as the number of samples increases) does not lead to substantial improvements in performance. These models consistently underperform relative to their coded counterparts, which are trained using the ELBO as the objective function. For a fair comparison, we use the same neural network architecture for the uncoded and the coded posterior (differing only in the final dense layer to match the respective output dimensionalities), the same network architecture for the decoder (differing only in the first dense layer to match the respective input dimensionalities), and the same simple independent prior.
Fig. 32 shows the evolution of the different objectives, with coded models obtaining better bounds than the uncoded IWAE models, even when considering 100 samples. In Table 9, we present test metrics, with coded models again overperforming their uncoded counterparts trained with the IWAE objective. We observe that some metrics slightly decline augmenting the number of samples, likely due to overfitting, since we applied a common early stopping point.
@@
Table 9: Comparison of test metrics between our method and the uncoded DVAE trained with the IWAE objective.
Model BER WER Entropy Acc. Conf. Acc. PSNR SSIM
uncoded 8 0.089 0.384 0.467 0.594 0.595 15.598 0.509
uncoded 8 IWAE 10 samples 0.063 0.372 1.309 0.617 0.640 14.282 0.470
uncoded 8 IWAE 20 samples 0.075 0.447 1.391 0.634 0.651 14.237 0.460
uncoded 8 IWAE 30 samples 0.074 0.438 1.564 0.619 0.641 13.757 0.457
uncoded 8 IWAE 100 samples 0.107 0.583 2.274 0.596 0.663 12.996 0.433
uncoded 8 IWAE ($\alpha=0.5$, 10 samples) 0.059 0.353 0.929 0.630 0.638 14.582 0.488
uncoded 8 IWAE ($\alpha=0.5$, 20 samples) 0.067 0.388 1.066 0.638 0.644 14.432 0.479
uncoded 8 IWAE ($\alpha=0.5$, 30 samples) 0.053 0.349 1.165 0.630 0.644 14.178 0.472
uncoded 8 IWAE (DR, $\alpha=0.5$, 10 samples) 0.052 0.312 1.004 0.623 0.629 14.542 0.483
uncoded 8 IWAE (DR, $\alpha=0.5$, 20 samples) 0.056 0.353 1.041 0.600 0.614 14.231 0.463
uncoded 8 IWAE (DR, $\alpha=0.5$, 30 samples) 0.057 0.352 1.068 0.602 0.607 14.016 0.449
coded 8/80 0.021 0.144 2.905 0.750 0.816 17.318 0.619
coded 8/160 0.027 0.189 3.637 0.783 0.831 17.713 0.641
coded 8/240 0.037 0.231 4.000 0.799 0.893 17.861 0.653
Appendix I Beyond Repetition Codes
  • Differentiability. Both encoding and decoding processes must be differentiable w.r.t. the inputs to enable gradient computation and backpropagation.
Since Shannon’s landmark work [Shannon, 1948], researchers have been developing feasible ECC schemes that meet (or approach) the aforementioned properties for ever-increasing values of $M$ and $R$. This is driven by the high throughput demands in digital communications and limited power constraints in the case of wireless communications. In this regard, state-of-the-art ECC schemes such as Low Density Parity Check (LDPC) codes [Gallager, 1962] can be designed to meet the above criteria, leveraging their linear algebraic structure. Similarly, encoding and decoding algorithms for polar codes [Arikan, 2009] have complexity of $\mathcal{O}(D\log D)$, approaching this target. Additionally, in digital communications, soft decision decoding is known to provide savings of up to 2-3 dB of signal energy over hard decision decoding (binary input). For this reason, significant effort has been also made in the communications field to develop efficient and powerful SISO decoders, such as the sum-product algorithm (SPA) [Kschischang et al., 2001] for LDPC codes.
Appendix J Hierarchical Coded-DVAE results
@@ Appendix I Beyond Repetition Codes
Inspired by polar codes [Arikan, 2009], we present a hierarchical Coded-DVAE with two layers of latent bits, as illustrated in Fig. 33. In this model, the latent bits $\boldsymbol{m}_{1}$ are encoded using a repetition code in the first layer, producing $\boldsymbol{c}_{1}$ and $\boldsymbol{z}_{1}$. Concurrently, the bits in the second layer, $\boldsymbol{m}_{2}$, are linearly combined with $\boldsymbol{m}_{1}$ following $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$, considering a binary field or Galois field. The resulting vector is then encoded with another repetition code to produce $\boldsymbol{c}_{2}$, which is subsequently modulated into $\boldsymbol{z}_{2}$. Finally, both $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ are concatenated and passed through the decoder network to generate $\boldsymbol{x}$. The model provides stronger protection for $\boldsymbol{m}_{1}$, as it appears in both branches of the generative model. Inference follows a similar approach to the one employed in the Coded-DVAE, incorporating the linear combination of $\boldsymbol{m}_{1}$ and $\boldsymbol{m}_{2}$ used in the second branch.
We adopt the same variational family as used in the standard Coded-VAE; however, in this case, we incorporate both hierarchical levels, leading to
@@
where $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ is calculated following the same approach as in the Coded-DVAE with repetition codes, computing the all-are-zero and all-are-ones products of probabilities of the bits in $\boldsymbol{c}_{1}$ that are copies of the same message bit. The posterior $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ considering both the encoder’s output and the inferred posterior distribution $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$. Note that in this case, the decoder outputs the probabilities for both $\boldsymbol{c}_{1}$ and $\boldsymbol{c}_{2}$, with $\boldsymbol{c}_{2}$ being the encoded version of the linear combination $\boldsymbol{m}_{1,2}=\boldsymbol{m}_{1}\oplus\boldsymbol{m}_{2}$. Consequently, we first obtain $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ following the same approach as in the Coded-DVAE with repetition codes, and determine $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})$ as
$q_{\boldsymbol{\eta}}(\boldsymbol{m}_{2}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(p_{j}),$ (14)
$p_{j}=q_{\boldsymbol{\eta}}(m_{1,2,j}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,j}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,2,j}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{1,j}=1|\boldsymbol{x}).$ (15)
@@
$q^{\prime}_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})=\prod_{j=1}^{M}\text{Ber}(q_{j}),$ (16)
$q_{j}=q_{\boldsymbol{\eta}}(m_{1,j}=1|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,j}=0|\boldsymbol{x})+q_{\boldsymbol{\eta}}(m_{1,j}=0|\boldsymbol{x})q_{\boldsymbol{\eta}}(m_{2,j}=1|\boldsymbol{x}).$ (17)
Next, we apply the soft encoding approach to incorporate the repetition code structure at both levels of the hierarchy. The posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1}|\boldsymbol{x})$ are repeated to obtain $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{1}|\boldsymbol{x})$, and the posterior probabilities $q_{\boldsymbol{\eta}}(\boldsymbol{m}_{1,2}|\boldsymbol{x})$ are repeated to produce $q_{\boldsymbol{\eta}}(\boldsymbol{c}_{2}|\boldsymbol{x})$. Utilizing the reparameterization trick from Eq. (5), we sample $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$, concatenate them to form $\boldsymbol{z}$, and pass this through the decoder to generate $p_{\boldsymbol{\theta}}(\boldsymbol{x}|\boldsymbol{z})$. The model is trained by maximizing the ELBO, given by
@@
We obtained results on the FMNIST dataset using a model with 5 information bits per branch and repetition rates of $R=1/10$ and $R=1/20$. In this case, we applied the same code rate to both branches, although varying code rates could be used to control the level of protection at each hierarchy level. Tables 10 and 11 present the metrics obtained for the different configurations. Specifically, Table 10 shows the overall metrics obtained with this structure, and Table 11 compares the error metrics across the two hierarchy levels. As expected, $\mathbf{m}_{2}$ shows poorer error metrics compared to $\mathbf{m}_{1}$, since the model provides more redundancy to $\mathbf{m}_{1}$ incorporating it in both branches. Although the overall metrics and generation quality are similar to those of the Coded-DVAE with 10 information bits (see tables in Figs. 5 and 4), the introduced hierarchy results in a more interpretable latent space. In this setup, $\mathbf{m}_{1}$ captures global features (such as clothing types in the FMNIST dataset), while $\mathbf{m}_{2}$ controls individual features, as we can observe in Fig. 34, where we show examples of the model’s generative outputs for fixed $\mathbf{m}_{1}$ and random samples of $\mathbf{m}_{2}$.
Figure 34: Examples of generated images using the hierarchical Coded-DVAE with (a) a 5/50 repetition code in each branch, and (b) a 5/100 repetition code in each branch. In all the examples provided, $\boldsymbol{m}_{1}$ was fixed while $\boldsymbol{m}_{2}$ was randomly sampled.
Table 10: Comparison of the obtained metrics for the Coded-DVAE with polar codes with different configurations, which we refer to as ‘hierarchical Coded-DVAE’.
@@
Appendix K Ablation study
@@
K.1 Ablation study on the hyperparameter $\beta$
@@
K.1.1 Coded model
We first evaluate the influence of the parameter $\beta$ in coded models. We take as a reference the coded model with 8 information bits and a rate $R=1/30$, and train it using $\beta=5,10,15,20$. We assess the performance of the model in reconstruction and
generation tasks. We observe the model is fairly robust, achieving similar performance across configurations in most metrics.
@@
K.2 Ablation study on the number of trainable parameters
A consistent architecture was employed across all experiments, which is detailed in Section C. However, since the introduction of the code alters the dimensionality of the latent space, it is necessary to adjust the encoder’s output and the decoder’s input. This results in an augmentation of the trainable parameters in the coded cases compared to their uncoded counterparts.
@@
Appendix L Evaluating log-likelihood using the soft-encoding model